OpenStudy (anonymous):

can someone help me? I have to turn in these assignments before 12 tonight and i have alot more stuff to do. ILL GIVE A MEDAL AND BECOME A FAN!

OpenStudy (anonymous):

One method you can use to determine whether a triangle is a right triangle, given three side lengths, is to apply the Converse of the Pythagorean Theorem. Alternately, you can use trigonometric ratios. Show that the triangle in the diagram is a right triangle by using trigonometric ratios. 

OpenStudy (anonymous):

@Luis_Rivera

jimthompson5910 (jim_thompson5910):

if you can show that a^2 + b^2 = c^2 is true, where c is the longest side, then you will have shown that the triangle is a right triangle

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 16^2 + 63^2 = 65^2 256 +3969 = 4225 4225 = 4225 so because the last equation is true, this means the first equation is true since a^2 + b^2 = c^2 is true when a = 16, b = 63 and c = 65, this means that the triangle is a right triangle

OpenStudy (anonymous):

You didn't have to give me the answer lol Leading me through it would have been helpful to but thank you so much! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

oh sry about that

jimthompson5910 (jim_thompson5910):

hopefully everything makes sense though?

OpenStudy (anonymous):

That's okay haha and yes it does i Had no clue how to do this what so ever.

jimthompson5910 (jim_thompson5910):

it's never come up in the lesson before?

jimthompson5910 (jim_thompson5910):

but that's good it's all clicking now

OpenStudy (anonymous):

It has but im home schooled and we learn everything in a chat room and i didn't make it to that class so I didn't have a clue what was going on

jimthompson5910 (jim_thompson5910):

I gotcha

OpenStudy (anonymous):

Could you help me with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

A surveyor is “shooting a line” to a point on a tree 55 m from his current position. After rotating his surveying instrument 25° to the left, he “shoots” another line to a point on a fence post 25 m away. a.) Determine the distance between the point on the tree to the point on the fence post. (Show the appropriate formula, substitutions, and work. Give the distance to the nearest hundredth of a meter.) b.) Use the Law of Sines to find the measure of T. (Show the appropriate formula, substitutions, and work. Give the measure of T to the nearest degree.) c.) Find the mF. 

jimthompson5910 (jim_thompson5910):

for part a), you need to use the law of cosines c^2 = a^2 + b^2 - 2ab*cos(C) c^2 = 25^2+55^2 - 2*25*55*cos(25) keep going to solve for c

jimthompson5910 (jim_thompson5910):

make sure you're in degree mode on your calculator